### v1

# Education-Oriented Graph Retrieval-Augmented Generation for Learning Path Recommendation

Abstract.
Learning path recommendation seeks to provide learners with a structured sequence of learning items (e.g., knowledge concepts or exercises) to optimize their learning efficiency. Despite significant efforts in this area, most existing methods primarily rely on prerequisite relationships, which present two major limitations: 1) Many educational datasets do not explicitly provide prerequisite relationships between knowledge concepts, hindering the application of current learning path recommendation methods. 2) Relying solely on prerequisite relationships as the sole knowledge structure can impede learning progress and negatively impact student outcomes. To address these challenges, we propose a novel approach, Discrimination Learning Enhances Learning Path Recommendation (DLELP), which enhances learning path recommendations by incorporating both prerequisite and similarity relationships between knowledge concepts. Specifically, we introduce a knowledge concept structure graph generation module that adaptively constructs knowledge concept structure graphs for different educational datasets, significantly improving the generalizability of learning path recommendation methods. We then propose a Discrimination Learning-driven Reinforcement Learning (DLRL) framework, which mitigates the issue of blocked learning paths, further enhancing the efficacy of learning path recommendations. Finally, we conduct extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides interpretable reasoning for the recommended learning paths.
Learning Path Recommendation, Discrimination Learning, Reinforcement Learning
copyright: noneconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: ccs: Applied computing E-learningccs: Information systems Recommender systems
1. Introduction
Figure 1. The comparison of traditional methods (top) and our proposed DLELP (bottom) is depicted. The top denotes traditional KC prerequisite relationship-based methods where confusion between similar KCs such as $\text{KC}_{3}$, $\text{KC}_{6}$, and $\text{KC}_{7}$ can hinder students’ mastery of $\text{KC}_{3}$ and impede their understanding of subsequent KCs. The bottom illustrates how DLELP enables students to more effectively differentiate between these similar KCs, thereby improving learning outcomes.
With the rapid advancement of online education platforms, the growing demand for personalized learning experiences has led to the increasing significance of learning path recommendation (LPR) (Wang et al., 2023; Zhang et al., 2021; Chen, 2009; Shi et al., 2020). It can adaptively recommend personalized learning paths based on individual learning tempo, enabling them to acquire knowledge and skills systematically and orderly, unlike the ”general education” approach in traditional education (Nabizadeh et al., 2017; Guan et al., 2025; Cheng et al., 2025).
Research on education has shown that the structure of the knowledge concept (KC) significantly influences the LPR (Liu et al., 2019). Existing solutions based on KC structure for LPR can be mainly divided into two parts: KC correlation-based, where personalized learning paths are provided to learners by leveraging the implicit correlations among knowledge concepts, and KC prerequisite relation-based, where personalized learning paths are offered to learners by leveraging the knowledge concepts prerequisite relationships. Among them, KC correlation-based methods, without considering the unambiguous KC structure, fail to precisely determine the customized learning tempo (Chen et al., 2023). In contrast, KC prerequisite relationship-based methods integrate learners’ knowledge mastery levels with the prerequisite relationships of learning materials to provide smooth and efficient learning paths (Liu et al., 2019; Li et al., 2023; Zhang et al., 2024; Li et al., 2024). Despite the significant success of these methods, they often encounter the following two challenges.
Most educational datasets do not explicitly provide prerequisite relationships between knowledge concepts, which hinders the application of existing LPR (Zhang et al., 2024). While few datasets provide corresponding KC prerequisite relationship graphs, these graphs often suffer from sparsity due to the incomplete knowledge concepts or inappropriate relationships in the proposed connections. Additionally, existing educational datasets rely on manually annotated KC prerequisite relationship graphs, which are labor-intensive to construct. Such issues restrict the applicability of existing learning path recommendation methods.
Relying solely on prerequisite relationships between KCs as the single knowledge structure can lead to ”blocked phenomena,” which hinders learning progress (Pan et al., 2024). A major cause is the lack of opportunities for learners to discriminate between similar KCs during the learning process. When confusion arises among these similar KCs, it impedes the mastery of key KC along the learning path, thereby obstructing the acquisition of target KC. According to Gagne’s learning hierarchy (Chinda, 2022), discrimination learning is a critical stage in the learning process, where learners develop the ability to distinguish or discriminate between two KCs (Rohrer, 2012). Figure 1 demonstrates that utilizing similar KCs for discrimination learning can significantly enhance the likelihood of a smoother learning path, improving both learning efficiency and outcomes. Existing traditional learning path recommendation methods lie in the exclusive reliance on sparse KC prerequisite relationships while disregarding KC similarity relationships, which consequently impedes the ability to recommend efficient learning paths that align with learners’ pacing.
To address the aforementioned challenges, this paper proposes a novel method named Discrimination Learning Enhances Learning Path Recommendation (DLELP) to achieve a more satisfactory learning path by considering dual KC structure prerequisite relationships and similarity relationships. Specifically, we propose a KC structure graph generation module to adaptively generate KC structure graphs. This module leverages TextGrad (Yuksekgonul et al., 2024) to mitigate hallucination issues in LLMs, thereby adaptively generating more accurate and densely connected KC structure graphs tailored to different educational datasets. Next, we develop a Discrimination Learning-driven Reinforcement Learning (DLRL) framework based on the generated KC prerequisite relationship graph and similarity relationship graph, which consists of a prerequisite agent and a similarity agent. The prerequisite agent determines whether learning should follow the prerequisite relationships among KCs. The similarity agent identifies easily confused KCs during blocked phenomena and integrates them into the learning path for discrimination learning.
Our main contributions are summarized as follows:
  • We propose a KC structure graph generation module that adaptively generates prerequisite relationship graphs and similarity relationship graphs for KCs based on different educational datasets, significantly enhancing the generalizability of various learning path recommendation methods.
  • We develop a discrimination learning-driven reinforcement learning (DLRL) framework that alleviates the blocked phenomena, further improving the performance of learning path recommendations.
  • We conduct extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides reasons for the recommended learning paths.
2. Related Work
2.1. Learning Path Recommendation
Learning path recommendation is a key task in online education, focused on crafting personalized, structured learning paths for different learners to optimize their learning outcomes. Existing methods can be categorized into two types according to the approach of KC structures considered: (1) KC correlation-based: learning paths are generated by inferring the latent correlations KCs. (2) KC prerequisite relationship-based: learning paths are generated based on the prerequisite relationships between KCs.
For branches of KC correlation-based approaches, many works have been proposed using different techniques, such as decision tree classifiers (Lin et al., 2013), matrix factorization (Nabizadeh et al., 2020), bayes theorem (Xu et al., 2012) etc. One of the key contributions is the work by Chen et al. (2023), who proposed the SRC framework. This framework models the complex relationships and interactions among KCs, enabling the generation of complete learning paths. Due to its ability to better account for the sequential dependencies between KCs, KC prerequisite relationships-based methods are rapidly gaining attention. This branch frequently treats learning path recommendation as a sequential decision-making problem, which results in the frequent application of advanced reinforcement learning techniques (Intayoad et al., 2020; Cai et al., 2018). For example, Liu et al. (2019) proposed a framework for adaptive learning utilizing cognitive structure, known as CSEAL, which integrates learners’ knowledge levels with the KC prerequisite relations of learning materials to provide personalized learning paths. Li et al. (2023) proposed a Graph Enhanced Hierarchical Reinforcement Learning (GEHRL) framework, which incorporates a graph-based candidate selector to constrain the action space of the low-level agent and employs a test-based internal reward mechanism to alleviate the issue of sparse external rewards. Zhang et al. (2024) proposed a difficulty driven hierarchical reinforcement learning (DLPR) framework, which constructed a hierarchical graph of learning and practice items to capture the difficulty of items and their higher-order relations. Despite their success, due to incomplete KC prerequisite relationships and neglect of KC similarity relationships, leading to inappropriate learning path recommendations.
2.2. Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) is a well-established method for answering user queries across large datasets(Ram et al., 2023; Laskar et al., 2020; Goodwin et al., 2020). It is specifically designed for scenarios where the answers are localized within specific text regions, and retrieving these regions provides adequate grounding for the subsequent generation task (Lewis et al., 2020). When applying large language models (LLMs), RAG first retrieves relevant information from external data sources and then incorporates this information into the context window of LLM along with the original query (Ram et al., 2023; Wang et al., 2024). Naïve RAG approaches achieve this by first converting documents into text, then splitting the text into chunks, and finally embedding these chunks into a vector space where similar positions correspond to similar semantics (Gao et al., 2023). In contrast to other RAG approaches, Edge et al. (2024) proposed a Graph RAG method that leverages global summarization of an LLM-derived knowledge graph, highlighting a previously unexplored aspect of graph structures in this context. However, Graph RAG methods typically require sufficiently accurate documentation for effective implementation. To obtain accurate documentation for KC descriptions, we integrated TextGrad (Yuksekgonul et al., 2024) and developed a KC structure graph generation module. This module mitigates the hallucination problem in LLMs and generates structural graphs of KCs based on their names, which not only adapts to generate structural graphs across different datasets containing KC names but also effectively supports most existing prerequisite-based learning path recommendation approaches.
3. Preliminaries
In this section, we discuss the formulation of learning path recommendations and the definition of Graph RAG (Edge et al., 2024).
3.1. Problem Definition
The primary objective of this paper is to solve the knowledge concept structure-based goal-oriented learning path recommendation problem. Suppose we have a set of exercise $E=\{e_{1},e_{2},\cdots,e_{M}\}$ and a set of knowledge concepts $C=\{c_{1},c_{2},\cdots,c_{N}\}$. Normally, for each learner, learning goals are denoted as $G=\{g_{1},g_{2},\cdots\}$, which is a subset of the exercise set $G\subset C$. Taking historical learning records $H=\{h_{0},h_{1},\cdots\,h_{i}\}$ as input and guided by the learning goal, a learning path $P=\{e_{1},e_{2},\cdots,e_{k}\}$ is recommended for learner. Each record $h_{i}=(e,score)$ contains a exercise $e$ and the corresponding performance score.
At the start of each learning session, learner gets an initial score $score_{i}$, where a score of $1$ indicates that the learner has already mastered the learning goal and 00 indicates otherwise. During each learning step, learner tackles an exercise $e_{t}$ and gives a feedback $\text{score}_{t}$. The series of exercises engaged by the learner constitutes the learning path. At the end of the session, a final feedback $\text{score}_{m}$ is provided by learning. To evaluate the effectiveness of the recommended learning path, we calculate the improvement in the learner’s feedback scores between the beginning and the end of the session. The effectiveness of the learning path recommendation, denoted as $E_{p}$ is computed as follows:
(1) $E_{p}=\frac{E_{\text{end}}-E_{\text{start}}}{E_{\text{sup}}-E_{\text{start}}},$
where $E_{\text{sup}}$ represents the maximum achievable score for the learning goals, which corresponds to the total number of learning objectives. Our objective is to maximize the $E_{p}$ by provide an effective learning path.
Figure 2. DLELP Framework Overview. The KC structure graph generation module on the left begins by inputting KC names into TextGrad, which then generates textual descriptions. These descriptions are subsequently used by the EDU-graph RAG to construct KC structure graphs. On the right, the learning path generation module initiates with the P-Agent and S-Agent filtering KCs. Following this, the D-Agent selects exercises of appropriate difficulty. The module then utilizes an environment KT to dynamically update the knowledge state in each session. Finally, the module generates the learning path step by step.
Figure 3. Knowledge Concept Analysis Objective.
3.2. Graph Retrieval-Augmented Generation
Graph RAG (Edge et al., 2024) is an open-source method developed by Microsoft, which is based on global summarization of an LLM-derived knowledge graph. It excels in capturing semantic information, extracting entities and relationships from text, and constructing corresponding knowledge graphs.
The workflow of Graph RAG involves two primary processes: indexing and querying. During the indexing process, the input textual data is divided into chunks. For each text unit, entities, relationships, and claims are identified and extracted. Using this extracted information, a knowledge graph is constructed. The knowledge graph is then hierarchically clustered to generate multiple community-level summaries. In the querying process, the knowledge graph and the generated summaries from the indexing stage are leveraged for inference based on the requirements of specific query tasks. Additionally, Graph RAG’s prompts can be tailored to meet the needs of various tasks across different domains, ensuring its adaptability to a wide range of applications.
To achieve a holistic understanding of the textual descriptions of knowledge concepts and generate accurate KC structure graphs, this paper introduces an Edu-Graph RAG method based on Graph RAG. This approach effectively alleviates the issue of insufficient textual content in knowledge concept descriptions within datasets and mitigates hallucination problems in LLMs, enabling the generation of prerequisite relation and similarity relation graphs for knowledge concept global summarization.
3.3. Framework Overview
Figure 2 presents an overview of DLELP framework, which consists of two primary modules. The left KC structure graph generation module addresses the limitation that most educational datasets do not explicitly provide prerequisite relationships and similarity relationships between KCs. This is achieved by leveraging graph retrieval-augmented generation method (Graph RAG) tailored to the education domain. By analyaing the textual descriptions of KCs, the module identifies prerequisite and similarity relationships to construct KC structure graphs. The right DLRL framework operates interactively during each session. It employs relationship agents that integrate both prerequisite relationships and similarity relationships among KCs to dynamically generate the final learning path.
4. KC Structure Graph Generation
In this section, we provide a concise overview of the KC structure graph generation module. First, we incorporate the TextGrad (Yuksekgonul et al., 2024) technique to analyze KC names, facilitating a more comprehensive acquisition of explanatory texts. Building on this, we design an EDU-Graph RAG approach to automatically generate KC structure graphs. By leveraging the generated textual descriptions, this module effectively captures both prerequisite and similarity relationships, constructing KC structure graphs that serve as a critical foundation for recommending appropriate learning paths.
4.1. KC Textual Analysis and Generation
To effectively capture the rich semantic relationships between KCs, we first parse them to generate detailed explanatory texts as a foundation for further analysis. Given that existing datasets for learning path recommendation lack detailed explanatory texts for KCs, we employ LLMs (Zhao et al., 2023) to generate these explanations, significantly lowering the effort required for manual verification and explanation generation. While LLMs can quickly retrieve and generate explanations via online searches, they are prone to hallucinations, which can lead to inaccurate explanations of the KCs.
To enhance the accuracy of knowledge concept analysis, we utilize the TextGrad (Yuksekgonul et al., 2024) method to evaluate the initial explanations generated by LLMs and iteratively refine them based on the evaluation results. The TextGrad workflow is represented as follows:
[…]
Here, we focus on the task of KC analysis optimization via TextGrad. In KC analysis optimization, the goal is to improve the analysis of overall KCs by continuously updating and optimizing the explanatory text for each KC, utilizing LLMs (e.g., gpt-4o) to generate more accurate descriptions. In particular, we often have a computation graph like the following:
(3) $\text{Objective}=\text{LLM}(\text{Question}+\text{Analysis}+\text{Evaluation}\,\,\text{Instruction}),$
where the parameter we optimize is the Analysis, and the loss function is obtained by an evaluation of the analysis.
In this process, the knowledge concept name serves as the input. Through prompt engineering, the LLM is guided to act as an expert in the education domain. It generates an initial explanation for the knowledge concept, which is subsequently evaluated by another LLM. Based on the evaluation results, further refinements are made. This iterative cycle of generating and evaluating explanations continues until an accurate and reliable explanation is produced. For instance, the objective function for the KC analysis is shown in Figure 7.
4.2. Education-Oriented Graph RAG
To automatically generate a knowledge graph that captures the prerequisite and similarity relationships among KCs, we present EDU-Graph RAG, an education-oriented adaptation of Graph RAG (Edge et al., 2024). Graph RAG aims to extract structured data from unstructured text using LLM and build labeled knowledge graphs to support various application scenarios, such as dataset question generation and summary question answering. It provides fine-tuning commands that allow adjustments to its prompt content for specific domains. By inputting customized instructions, Graph RAG is tailored to specialize in analyzing the relationships between KCs in the education field. Moreover, prompt adjustments ensure it can accurately capture prerequisite relationships and similarity relationships among KCs, improving its domain-specific effectiveness. EDU-Graph RAG is tailored specifically for the education domain by fine-tuning its prompt to process the explanatory texts of KCs and extract prerequisite relationships and similarity relationships. The operational process of EDU-Graph RAG is outlined as follows:
Text Chunking. The explanatory texts of knowledge concepts are divided into smaller, manageable chunks suitable for LLM processing. This enhances both recall and precision in analyzing knowledge concept texts.
Extraction of Element Instances. Entities and associated statements are extracted from the parsed explanatory texts. For example, this includes the professional domain of the KC and its prerequisite relationships and similarity relationships with other KCs.
Generation of Element Summaries. Based on the extracted entities and statements, meaningful summaries are created. These summaries may include implied relationships and conceptual overv-
iews that are not explicitly stated in the text.
KC Structure Graph Construction. Finally, using the extracted entities, statements, and summaries, a comprehensive knowledge graph of prerequisite relationships and similarity relationships among KCs is constructed.
This process ensures an accurate and structured representation of prerequisite relationships and similarity relationships, providing the necessary KC structure graph for the learning path recommendation model, which in turn generates personalized learning paths based on learners’ needs and objectives.
5. Learning Path Generation
In this section, we develop a Discrimination Learning-driven Reinforcement Learning (DLRL) framework that alleviates the blocked phenomenon, further improving the performance of learning path recommendations. To recommend an effective learning path, we incorporate a Difficulty Matching Knowledge Tracing (DIMKT) model (Shen et al., 2022) as the environment KT model to dynamically track the evolving knowledge states of learners. Furthermore, we introduce a dual-agent approach that integrates a Prerequisite Agent (P-Agent) and a Similarity Agent (S-Agent), which work collaboratively to monitor the learner’s progress and adaptively select KCs based on either prerequisite or similarity relationships. And a Difficulty Agent (D-Agent), which operates by analyzing the selected KCs and the learner’s mastery level of these KCs. It selects exercises that align with the learner’s current difficulty level while maintaining a consistent and gradual progression in the difficulty of the recommended learning path. This ensures that the learning path remains smooth and appropriately challenging, maximizing the learner’s engagement and effectiveness.
5.1. Knowledge Tracing
The dynamic changes in a learner’s knowledge mastery level during the learning process serve as a reflection of their learning outcomes. By leveraging these insights to devise subsequent recommendation strategies, it is possible to provide a learning path that aligns with the learner’s abilities and progress. Given the maturity and robustness of existing knowledge tracing (KT) methods, such as the PYKT knowledge tracing library (Liu et al., 2022), which allows for the selection of suitable KT approaches to predict learners’ knowledge states, we adopt the DIMKT method for tracking learners’ knowledge mastery level dynamics. DIMKT is particularly advantageous as it can predict the impact of exercise difficulty on learners, thereby enhancing the accuracy of recommending exercises based on their difficulty level.
In this process, we employ the same HGNN (Yang et al., 2023; Zhang et al., 2023; Zheng et al., 2023) embedding technique as utilized in DLPR (Zhang et al., 2024) to embed both exercises and difficulty levels as inputs to DIMKT. This method effectively integrates exercise features with difficulty vectors, ensuring that the predicted knowledge states comprehensively account for the interrelations between exercises and their corresponding difficulty levels.
5.2. Prerequisite Agent
The P-agent is designed to recommend a suitable learning path for learners based on KCs prerequisite relationships. However, due to the vast number of knowledge concepts, this selection process results in a large search space. To reduce this complexity, it is essential to identify a preliminary shortest learning path between an initial node and the learning objective before the P-agent performs its selection. Existing advanced methods, such as adaptive action space learning algorithms, employ the A* algorithm (Duchoň et al., 2014) to generate the shortest paths while dynamically determining candidate action spaces. Nonetheless, these methods often fail to identify an effective initial node, which may result in the selection of unsuitable knowledge concepts at the outset and ultimately lead to an ineffective learning path.
To address this limitation, we introduce an initial node identification algorithm to ensure the selection of the most appropriate initial node. Specifically, our approach begins from the target knowledge concept node and iteratively traces back along the prerequisite relationship. During this backward traversal, we assess whether the prerequisite knowledge concepts of the current node have already been mastered by the learner. If these prerequisite concepts are mastered, the backward traversal terminates. Otherwise, we continue tracing back to the node with the longest chain of prerequisite knowledge concepts. By incorporating this method, the most suitable initial node can be determined before dynamically identifying the candidate action space, effectively optimizing the overall learning path selection process.
5.2.1. State encoder.
At each step, the predicted knowledge state is based on the learner’s learning objectives and historical learning records. Thus, at the end of the $t-1$ step, the state of the P-agent includes the learner’s knowledge state $\mathbf{h}_{t-1}$ and learning objectives $\mathbf{G}$. Specifically, we use a one-hot encoding scheme to represent the learning objective $\mathbf{G}=\{0,1\}^{M}$, where $M$ is the total number of knowledge concepts. The index corresponding to the learning objective is set to $1$, while all other indices are set to 00. At step $t$, the state of the P-agent $\mathbf{s}_{t}$ is represented as follows:
(4) $\mathbf{s}_{t}=\mathbf{h}_{t-1}\oplus\mathbf{G}.$
5.2.2. Policy.
We adopt Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the model for the prerequisite relation agent, as it is one of the most performant reinforcement learning models currently available. PPO is particularly effective at identifying the optimal sequence of knowledge concepts to achieve the learning goal. Specifically, PPO consists of two primary components: a policy network (Actor) $\pi(\mathbf{G}\left|\mathbf{s}_{t};\theta\right.)$ and a value network (Critic) $V(\mathbf{s}_{t};\phi)$, where $\theta$ and $\phi$ represent the respective network parameters.
The goal of the policy network is to generate a probability distribution over the knowledge concepts in the dataset. The state of the P-agent is fed into the policy network, which outputs a corresponding probability distribution over the knowledge concepts.
(5) $\mathbf{G}\sim\pi(\mathbf{G}\left|\mathbf{s}_{t};\theta\right.)=\text{Softmax}(\text{FC}(\mathbf{s}_{t})),$
where $FC$ is the fully connected layer. The value network takes the state of the relation agent as input and evaluates the reward associated with that state.
(6) $V\left(\mathbf{s}_{t};\phi\right)=\text{FC}\left(\mathbf{s}_{t}\right).$
During each training step, we optimize PPO using a loss function similar to GEHRL (Chen et al., 2023). The value network is trained using the mean squared error (MSE) loss function with gradient descent.
(7) $L\left(\phi\right)=\mathbb{E}\left(\left\|\sum_{i=0}^{T-t}{\gamma^{i}Re_{t+i}-V\left(\mathbf{s}_{t};\phi\right)}\right\|\right)^{2},$
where $Re_{t}$ represents the reward at step $t$, the computation of which will be detailed later, and $\gamma$ denotes the discount factor. Based on existing reinforcement learning optimization algorithms for sequential decision-making, we optimized the policy network using PPO-clip (Chen et al., 2023).
(8) $\hat{A}_{t}=-V\left(\mathbf{s}_{t};\phi\right)+Re_{t}+\cdots+\gamma^{T-t+1}Re_{T-1}+\gamma^{T-t}V\left(\mathbf{s}_{T};\phi\right),$
(9) $L(\theta)=\hat{\mathbb{E}}_{t}\left[\text{min}\left(r_{t}\left(\theta\right)\hat{A}_{t},clip\left(r_{t}\left(\theta\right),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right],$
where $r_{t}(\theta)$ represents the probability ratio $\frac{\pi(\left.G\right|s_{t};\theta)}{\pi\left(\left.G\right|s_{t};\theta_{old}\right)}$, $\epsilon=0.2$ is a hyperparameter, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}$ modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $r_{t}$ outside of the interval $[1-\epsilon,1+\epsilon]$.
5.2.3. Reward
To maximize the improvement in learners’ mastery of the learning objectives, we calculate the corresponding reward value after selecting knowledge concepts at each step, as described in Section 3.1.
(10) $Re_{t}=\begin{cases}E_{p},\,\,\text{if t is the last learning stage}\\ 0,\,\,\,\,\text{otherwise.}\\ \end{cases}$
5.3. Similarity Agent
While the prerequisite relationships between knowledge concepts provide a basis for recommending a learning path, they often exhibit limited efficacy in significantly enhancing a learner’s understanding of the current knowledge concept. To address this issue, we introduced a similarity agent (S-agent). The S-agent aims to identify knowledge concepts that are closely related to the target knowledge concept. These similar knowledge concepts can effectively support the learner’s comprehension by reinforcing their grasp of the target knowledge concept. By leveraging the similarity relationship, the S-agent ensures that the recommended learning path not only adheres to prerequisite relationships but also optimally promotes the learner’s mastery of the target knowledge concepts.
At step $t$, the P-agent evaluates the improvement in the learner’s mastery of prerequisite knowledge concept $c_{t-1}$ when selecting the next knowledge concept $c_{t}$ to recommend. The evaluation is based on the following:
(11) $p_{t}=h_{t-1}^{c_{t}}-h_{t-2}^{c_{t}},$
where $h_{t-1}^{c_{t}}$ represents the learner’s knowledge mastery level associated with the knowledge concept $c_{t}$ learned at step $t-1$, and $h_{t-2}^{c_{t}}$ represents the knowledge mastery level at step $t-2$.
If the improvement $p_{t}$ is below a predefined threshold $\tau$, the S-agent is activated. The S-agent identifies a set of similar knowledge concepts that are related to the current knowledge concept $c_{t}$. From this set, the S-agent selects a sub-path that can effectively enhance the learner’s mastery of the current knowledge concept. This sub-path is then appended to the learning path generated by the P-agent. Once the S-agent completes its selecting process, the P-agent resumes its operation, selecting subsequent knowledge concepts based on their prerequisite relationship.
Table 1. Performance comparison for learning path recommendation methods. Existing state-of-the-art results are underlined and the best results are bold. Our DLELP is compared with the SOTA DLPR and * indicates a $p-value<0.05$ in the t-test.
Dataset Steps KNN GRU4Rec RL-Tutor Actor-Critic CSEAL SRC GEHRL DLPR DLELP
Junyi 5 0.1296 0.1504 0.1601 0.1891 0.1964 0.1874 0.1762 0.2086 0.2406
10 0.1485 0.2129 0.1923 0.2051 0.1779 0.2075 0.2222 0.2293 0.2431
15 0.1769 0.1714 0.1272 0.1428 0.1981 0.1368 0.2105 0.2181 0.2295
20 0.1415 0.1909 0.1052 0.2129 0.1880 0.1739 0.2358 0.1880 0.2758
MOOCCubeX 5 0.2131 0.2500 0.2500 0.2501 0.2686 0.2131 0.1164 0.2388 0.3194
10 0.1304 0.1846 0.2001 0.2089 0.1718 0.2203 0.1342 0.2173 0.3508
15 0.1791 0.2343 0.1875 0.1818 0.2328 0.2695 0.1689 0.2933 0.3205
20 0.2388 0.2272 0.1600 0.2352 0.2676 0.2631 0.1063 0.2781 0.3589
ASSIST09 5 0.0504 0.0771 0.0625 0.0498 0.0971 0.0759 0.0966 0.1245 0.1268
10 0.0916 0.0849 0.0750 0.0696 0.1024 0.0760 0.1018 0.0926 0.1174
15 0.0734 0.0482 0.0843 0.0836 0.0931 0.1273 0.1183 0.0996 0.1281
20 0.0666 0.0675 0.1111 0.0616 0.0769 0.1417 0.1195 0.1434 0.1455
5.4. Difficulty Agent
After the P-Agent and S-Agent select the knowledge concepts, we incorporate a difficulty agent to match the selected concepts with exercises that align with the learner’s knowledge mastery level. This approach prevents the recommended exercises from being too difficult or too easy, which could lead to a loss of interest during the learning process (Huang et al., 2017; Minn et al., 2018; Zhang et al., 2022). Each knowledge concept may correspond to one or more exercises with varying levels of difficulty. For knowledge concepts where the learner’s mastery is low, easier exercises are required. Conversely, for concepts where mastery is high, more challenging exercises are necessary. Therefore, we need to find the exercise $e_{m}$ whose difficulty is closest to the learner’s knowledge state.
(12) $m=\text{argmin}_{m}\left|\text{difficult}\left(e_{m}^{c_{i}}\right)-h^{c_{i}}\right|,$
[…]
6. Experiments
In this section, we first introduce the experimental setup. Then, through extensive experiments, we demonstrate the superiority of our method based on the experimental results.
6.1. Datasets
We conduct our experiments using three publicly available educational datasets to validate the effectiveness of our model: Junyi  [^1] https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1198, MOOCCubeX  [^2] https://github.com/THU-KEG/MOOCCubeX - Computer, and ASSISTments2009  [^3] https://sites.google.com/site/assistmentsdata/home/2009-2010-assistment-data (ASSIST09). All three educational datasets include learner activity data, which comprises learning records as well as the corresponding knowledge concept names. Please note that the Junyi dataset provides a comprehensive KC structure graph, illustrating the prerequisite relationships among knowledge concepts. In contrast, the MOOCCubex dataset contains an incomplete KC structure graph, while the ASSIST09 dataset lacks a KC structure graph altogether. Furthermore, none of the datasets include information on the similarity relationships between KCs. However, all three datasets contain fields with knowledge concept names, allowing us to construct KC structure graphs for both prerequisite and similarity relationships using our method. The statistical details of the datasets are shown in Table 2.
Table 2. Dataset Statistics
Dataset Junyi MOOCCubeX ASSIST09
#KCs 835 443 167
#Learners 525,061 629 4,217
#Records 21,460,249 17,447 346,860
Positive label rate 54.38% 70.18% 63.81%
6.2. Compared Methods
To demonstrate the effectiveness of our method, we compare it with the following baseline methods:
  • KNN (Cover and Hart, 1967): The K-Nearest Neighbors algorithm identifies the K closest neighboring values to the target. We use the cosine similarity between learning paths to measure the similarity between two learners and recommend the most similar learning paths.
  • GRU4Rec (Hidasi, 2015): GRU4Rec is a session-based recommendation model based on Gated Recurrent Units (GRU). It trains an RNN model to capture patterns and information within sequences, predicting the probability of the next exercise appearing.
  • Actor-Critic (Konda and Tsitsiklis, 1999): This method encodes input data using a GRU and then employs a vanilla actor-critic framework for learning path recommendation.
  • RL-Tutor (Kubotani et al., 2021): RL-Tutor is an RL-based adaptive tutoring system that simulates student behavior using DAS3H (Dwivedi et al., 2018) and recommends learning paths via reinforcement learning.
  • CSEAL (Liu et al., 2019): CSEAL uses DKT (Piech et al., 2015) to estimate knowledge states and designs a cognitive navigation algorithm to guide the vanilla actor-critic framework in recommending learning paths.
  • SRC (Chen et al., 2023): SRC uses a concept-aware encoder to capture the relationships between knowledge concepts, and then a decoder generates learning paths.
  • GEHRL (Li et al., 2023): GEHRL employs a hierarchical reinforcement learning approach, where the lower-level agent filters relevant knowledge concepts for the target, while the higher-level agent is responsible for planning the sequence.
  • DLPR (Zhang et al., 2024): DLPR uses a hierarchical graph neural network to aggregate information about learning tasks and difficulty and then applies a hierarchical reinforcement learning framework to plan knowledge concepts and recommend exercises.
We evaluate these methods based on the overall improvement of learning objectives after recommendation. Specifically, in Eq. (1), $E_{\text{sup}}$ represents the total sum of goals for all learners, while $E_{\text{start}}$ and $E_{\text{end}}$ represent the total sum of initial and final scores for all learners, respectively.
6.3. Implementation Detail
We implemented our proposed learning path recommendation framework using PyTorch (Paszke et al., 2019) and the simulator code from previous works(Li et al., 2023; Zhang et al., 2024). In the process of generating KC structures graph, we adjusted the value of the chunk based on the number of KCs in the dataset. The chunk values used for the Junyi, Assist09, and MOOCCubex datasets were $800$, $600$, and $400$, respectively. In the process of recommending learning path, the first 60% of exercise logs in the dataset were used to train the environment KT, while the remaining exercise logs were used to train and test our model. To more accurately evaluate the model’s performance, we trained a DKT (Piech et al., 2015) on the entire dataset to predict learners’ mastery of target knowledge concepts in Eq. (1). Further, similar to previous methods (Zhang et al., 2024; Li et al., 2023; Liu et al., 2019), we considered a learner’s mastery of a KC to be greater than $0.5$, as this implies a probability greater than $0.5$ of correctly answering an exericise related to that KC. Through testing, we observed that when using our DKT (Piech et al., 2015) for evaluation, the average improvement in learners’ knowledge state per exercise was $0.001$. Thus, in section 5.3, we set the threshold $\tau$ to 0.001. We used the Adam optimizer (Diederik and Ba, 2015) for both the policy network and the value network, with a learning rate of 0.001. Our code is available at https://anonymous.4open.science/r/DLELP.
6.4. Overall Performance
Table 1 presents the overall performance of all methods across the three datasets, revealing several key insights. In comparison with all baseline models, our method outperforms the others on all three datasets. Specifically, among all learning path recommendation methods, our approach surpasses KC correlation-based methods (Chen et al., 2023), as well as KC prerequisite relation-based methods that consider knowledge concept structures and changes in learners’ knowledge states during sequential decision-making, such as CSEAL (Liu et al., 2019), GEHRL (Li et al., 2023), and DLPR (Zhang et al., 2024). These advanced prerequisite relation-based methods are outperformed by our approach, demonstrating the progress our method makes compared to previous methods. Furthermore, in each dataset, performance generally fluctuates with the variation in recommended steps. This indicates that, aside from the fact that recommending different exercises can have varying effects on learners, the number of recommended exercises also plays a role. When step=20, the recommended learning paths achieved optimal effects across all three datasets. This indicates that the learning path of this length is relatively suitable for the overall learners, enabling them to better grasp the target knowledge concept during the learning process.
Table 3. Ablation Experiment: ”w/o S” represents the model without similarity agent, and ”w S” represents the overall model.
Junyi MOOCCubeX ASSIST09
w S w/o S w S w/o S w S w/o S
step=5 0.2406 0.1932 0.3194 0.2148 0.1268 0.0737
step=10 0.2431 0.2051 0.3508 0.1928 0.1174 0.1007
step=15 0.2295 0.1851 0.3205 0.2258 0.1281 0.1190
step=15 0.2758 0.2232 0.3589 0.1517 0.1455 0.1264
6.5. Ablation Study
To validate the effectiveness of the innovative components of our method, we conducted an ablation study. Specifically, we removed the similarity agent module from our method to assess its contribution to the overall performance. As shown in Table LABEL:tab:Ablation_Experiment, when the similarity agent is removed from our method, there is a noticeable performance drop across all three datasets. This highlights the necessity and value of considering the similarity relationships between knowledge concepts in learning path recommendations.
Figure 4. Comparison between the original graph and the graph generated by DLELP. KC denotes the number of knowledge concepts, while Relationship represents the number of prerequisite relationships among these KCs.
Figure 5. Performance comparison of the original graph and the DLELP-generated graph.
6.6. Knowledge Graph Generation
To demonstrate the effectiveness of the KC prerequisite relationship graph generated by our method, we visualized both the original and generated knowledge graphs for the Junyi and MOOCCubeX datasets. The visualization results of the KC prerequisite relationship graphs are shown in Figure 4. It is important to note that since ASSIST09 does not explicitly provide the KC prerequisite relationship graph, it is not included in the analysis. We observe that in the Junyi dataset, the number of relationships in the generated KC prerequisite relationship graph is higher than in the original graph, indicating that our method captures more and more complex prerequisite relationships between KCs. In the MOOCCubeX dataset, the number of KCs in the generated KC prerequisite relationship graph is greater than that in the original graph, enabling the inclusion of a more comprehensive set of KCs. This helps avoid the issue of insufficient KCs in the graph, which could otherwise limit the applicability of the model. Notably, the number of relationships in MOOCCubeX indicates that each KC is connected to every other KC, which is evidently incorrect for the original graph.
Furthermore, we compared the performance of models using the original KC prerequisite relationship graph and the graph generated by our method on the Junyi dataset. Please note that similarity relationship graphs are uniformly generated by DLELP. From Figure 5, we find that the performance of the model using the generated KC prerequisite relationship graph is comparable to that using the original KC prerequisite relationship graph. This demonstrates that, in addition to automatically generating the prerequisite relationship graph, our method can produce KC prerequisite relationship graphs that perform at a similar level to the original graph, which was manually constructed by experts based on domain knowledge and a deep understanding of the relationships between KCs.
Figure 6. The explanation of learning path.
Figure 7. Compare the Results of with/without TextGrad.
6.7. Case Study
Figure 7 shows the explanation outputs of the KCs by the LLM (gpt-4o) before and after using TextGrad. When explaining the knowledge concepts ”logical address” and ”virtual address” using the LLM alone, confusion arises, as these two KCs are easily mixed up. The LLM would incorrectly explain that the ”logical address” is the same as the ”virtual address” . After optimizing the explanation process using the TextGrad method, it is clear that the LLM can effectively distinguish between the two KCs.
In the process of generating the KC structure graph, EDU-Graph RAG also generates a community summary containing key entities, relationships, and claims, which provides useful contextual information for subsequent queries. Taking advantage of this feature of EDU-Graph RAG, we analyzed and explained the recommended learning paths, which improves the interpretability of the recommendations and enhances learners’ trust in the recommended content. Figure 6 shows the analysis and explanation of the recommended paths by EDU-Graph RAG. For each recommended exercise, EDU-Graph RAG is able to explain the reasoning behind the recommendation of the corresponding KC.
Figure 8. Results of simulation experiment.
6.8. Simulation Experiment
Since real-world data only contains static information, which cannot be directly used to analyze the exercise sequences of students not included in the dataset, the effectiveness of our method in real online education scenarios cannot be directly validated. Therefore, we used a Knowledge Evolution-based Simulator (KES), which leverages the DKT model (Piech et al., 2015) to simulate students’ exercise behavior on randomly generated exercise sequences. Subsequently, our method performs learning path recommendations based on the student behavior generated by the simulator to validate its effectiveness in real online education scenarios.
We constructed the KES-Junyi simulator (Liu et al., 2019) based on the Junyi dataset. Using simulated initial data, we compared our method with three advanced learning path recommendation methods—SRC, GEHRL, and DLPR—to demonstrate the effectiveness of our approach in online education settings. The results of the simulation experiments are shown in Figure 8. We observe that our method performs better than the other models on the simulated data. This indicates that, when faced with complex and dynamic learner information, our method exhibits strong adaptability, outperforming existing advanced learning path recommendation methods, and is well-suited for application in online education scenarios.
7. Conclusion
In this paper, we proposed a DLELP method to achieve a more effective learning path by considering dual KC structures prerequisite relationships and similarity relationships. To be specific, we proposed a knowledge concept structure graph generation module that adaptively generates graphs for knowledge concepts based on diverse educational datasets, enhancing the generalizability of existing learning path recommendation methods. Then, we developed a Discrimination Learning-driven Reinforcement Learning (DLRL) framework that alleviates the blocked phenomenon, further improving the performance of learning path recommendations. Finally, we conducted extensive experiments on three benchmark datasets, demonstrating that our method not only achieves state-of-the-art performance but also provides reasons for the recommended learning paths.